OpenStudy (anonymous):

In graphing y = -5(x+2)2 + 1, the vertex of the parabola is at what point

OpenStudy (amriju):

write this in the form of x^2=-4ay...so that would be lyk (x+2)^2=-4.(1/20)(y-1)...u see what I did there??..now ur vertex is at x+2=0 ie x=-2...and y-1=0...that gives y=1...so (-2,1) is the point...get it??..its done lyk dis

OpenStudy (anonymous):

omg I get it Thanks

OpenStudy (anonymous):

can you help me with one more question please ?

OpenStudy (amriju):

m sry i went offline yestrdy